OpenStudy (anonymous):

-2x+4y=8 solve for y

OpenStudy (akashdeepdeb):

Whenever you have these kinds of expressions You will get the answer in terms of the other variable. So what you've got to do is to keep the y on one side [RHS preferrably] And then Take ALL OTHER TERMS TO the LHS. Getting me at this? :)

OpenStudy (akashdeepdeb):

-2x+4y=8 4y = 8 + 2x [ADDING 2x to both sides to get rid of it in RHS] And now we have to get rid of the 4 in the RHS. So divide BOTH sides by 4 for it to REMAIN equal 4y/4 = (2x + 8)/4 y = x/2 + 2 Understood? :)

OpenStudy (anonymous):

I think so

OpenStudy (akashdeepdeb):

Let us take another example x + 2y - 3 = 0 And now we have to get x here [So, now we know that we'd get it in terms of y] x = +3 - 2y This is EXACTLY like solving a Simple equation! :D Got it? :)

OpenStudy (anonymous):

thank you yes I understand

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

can you help with one more

OpenStudy (anonymous):

When solving for such equations, you basically do the opposite of PEMDAS, taking care of the AS (addition/subtraction) first, then the MD (multiplication/division) so if we were to do so with this equation: \[-2x+4y=8\]what we are trying to do is get \(y\) by itself take away all the extra terms from around the \(y\) term:\[-2x+2x+4y=8+2x\]\[4y=8+2x\]we still have a \(4\) tagged to the \(y\) and since it's being multiplied we would simply divide to remove it:\[{4y\over4}={8\over4}+{2x\over4}\]\[y=2+{1\over2}x\]

OpenStudy (preetha):

Yummydum - impresssive!

OpenStudy (anonymous):

Thanks, @Preetha

OpenStudy (anonymous):

can some help with this one as well 6x-3y=-12 solve for y

OpenStudy (anonymous):

you do the same, opposite of PEMDAS:\[6x-3y=-12\]get rid of everything around the \(y\) term first :) what do you get?

OpenStudy (akashdeepdeb):

Firstly you can divide the whole equation by 3 Just for the sake of simplicity So it becomes 2x - y = -4 Now it is very simple to find y Take y to a preferable side [LHS] 2x + 4 = y OR y = 2x + 4 SOLVE ALL THESE TYPES QUESTIONS BY THE SIMILAR METHODS!! :D Understood? :)

OpenStudy (anonymous):

will it be 3=12+6x

OpenStudy (anonymous):

I meant 3y=-12+6x

OpenStudy (akashdeepdeb):

Yes 3y = 12 + 6x And then to remove the 3 from the y We'd have to DIVIDE the whole equation by 3 To get the final answer of y, interms of x.

OpenStudy (anonymous):

yes and then you would divide it all by \(3\) to be left with just \(y\) :)

OpenStudy (akashdeepdeb):

So Long....

OpenStudy (anonymous):

What's so long? @AkashdeepDeb

OpenStudy (akashdeepdeb):

It is a way of saying Bye (I guess). So Long and thanks for all the fish.... :D

OpenStudy (anonymous):

oh lol